# SHIELD: Secure Hypernetworks for Incremental Expansion Learning Defense

### v2

@@ Appendix H Ablation Study
Figure 7: AA of SHIELD and HNET under PGD attacks with increasing number of gradient steps, evaluated on models trained on Permuted MNIST (Figure 7(a)), Split CIFAR-100 (Figure 7(b)), and Split miniImageNet (Figure 7(c)).
Impact of $\beta$ Parameter Selection on Training Process of SHIELD
We evaluate the ability of SHIELD to retain knowledge across tasks, which is controlled by the hyperparameter $\beta$. In Figure 8, we present AA curves obtained by training the model with optimal hyperparameters while varying $\beta$. The results show that $\beta$ values of 0.010.010.01 and 1.01.01.0 yield the best performance, ensuring stable predictions and minimal forgetting. In contrast, a small value of $\beta=0.001$ leads to significant forgetting, particularly on the 5th and 6th tasks, after learning the final task.
@@
Figure 8: Comparison of AA curves for varying $\beta$ values, which control hypernetwork regularization strength in SHIELD
Performance of SHIELD on TinyImageNet dataset
We assess the scalability of SHIELD on a more challenging benchmark involving a larger number of tasks. Specifically, we use TinyImageNet, a subset of ImageNet, consisting of 200 classes, each with 500 training images and 50 validation images at a resolution of $\times 64$ 64 pixels. To construct the benchmark, we define 40 tasks, each composed of 5 randomly selected, non-overlapping classes. This setup allows us to evaluate SHIELD’s performance under increased task complexity and class diversity.
We compare the robustness of SHIELD with the traditional HNET approach (von Oswald et al. 2019) under three adversarial attacks: FGSM, PGD, and AutoAttack. Both models are trained specifically on TinyImageNet. For FGSM, we set the attack strength to $\varepsilon_{\text{attack}}=\frac{4}{255}$. For PGD and AutoAttack, we use $\varepsilon_{\text{attack}}=\frac{2}{255}$, with a PGD step size of $\delta=\frac{4}{255}$ and 100 iterations. The AutoAttack configuration is the same as used in the main experiments of the paper.
@@
Table 9: Comparison of AA after completing all tasks on the TinyImageNet dataset.
@@
From Table 9, we observe that even on a challenging dataset like TinyImageNet, our model achieves high AA on original (non-adversarial) test samples while also maintaining robustness against adversarial attacks. While HNET achieves competitive results, our method performs slightly better, especially under adversarial conditions.
Appendix I Extended Results from Main Paper
In this section, we extend results from Table 1 to attach standard deviations. We present extended results in Table 10, 11, 12, and 13.
Table 10: Comparisons of AA and BWT (only for original samples) after learning all the tasks on Permuted MNIST dataset. The results for baselines were extracted from Table 1 in (Ru et al. 2024).
@@ Appendix I Extended Results from Main Paper
SHIELDIM 80.08$\pm$0.01 97.44 $\pm$ 0.05 79.09 $\pm$ 0.87 97.96 $\pm$ 0.04 -0.05 $\pm$ 0.03
Table 11: Comparisons of AA and BWT (only for original samples) after learning all the tasks on Rotated MNIST dataset. The results for baselines were extracted from Table 2 in (Ru et al. 2024).
@@
Table 12: Comparisons of AA and BWT (only for original samples) after learning all the tasks on Split CIFAR-100 dataset. The results for baselines were extracted from Table 3 in (Ru et al. 2024).
@@
SHIELDIM 63.08 $\pm$ 0.64 62.39$\pm$ 0.38 46.48 $\pm$ 0.33 67.45 $\pm$ 0.28 -0.41 $\pm$ 0.19
  • Note: In the original Table 3 from (Ru et al. 2024), the reported accuracy for FGSM was higher than for original samples, which is inconsistent with expected adversarial behavior. We assume this was a mistake and have swapped the columns accordingly for a fair comparison.
Table 13: Comparisons of AA and BWT (only for original samples) after learning all the tasks on Split miniImageNet dataset. The results for baselines were extracted from Figure 3 in (Ru et al. 2024).
@@
Appendix J Class-Incremental Learning Results
Table 14 presents a comparison between SHIELD and the HNET baseline in terms of adversarial robustness across four datasets: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet, within the Class-Incremental Learning (CIL) setting. The results include AA under three adversarial attacks (AutoAttack, PGD, FGSM) as well as accuracy on original test samples, all reported after learning the full task sequence.
@@
The results reported in Table 14 were averaged over multiple random seeds to ensure statistical reliability. Specifically, results for Permuted MNIST, Rotated MNIST, and Split CIFAR-100 were averaged over 5 seeds, while Split miniImageNet results were averaged over 3 seeds. Due to the computational cost of AutoAttack, its results were averaged over 2 seeds for all datasets.
Table 14: Adversarial robustness of SHIELD across multiple datasets in the CIL scenario. We report AA under AutoAttack, PGD, and FGSM attacks, as well as on original samples after learning all tasks.
@@
[…]
Appendix K Training Algorithm
@@
Appendix L Training Time of SHIELD
Table 15 reports the training time of SHIELD model across different datasets. The durations are shown in the HH:MM:SS format and include standard deviations where available. For Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet, results are averaged over 5 training runs using the best-performing hyperparameters. For TinyImageNet, the time is reported from a single run with no standard deviation.
Table 15: Comparison of training time of the SHIELD model on considered datasets.
@@
Permuted MNIST 00:37:59±00:00:0200:37:59$\pm$ 00:00:0200 : 37 : 59 ± 00 : 00 : 02
Rotated MNIST 00:36:18±00:00:1900:36:18$\pm$ 00:00:1900 : 36 : 18 ± 00 : 00 : 19
Split CIFAR-100 02:10:23±00:02:0002:10:23$\pm$ 00:02:0002 : 10 : 23 ± 00 : 02 : 00
Split miniImageNet 09:21:05±00:12:1109:21:05$\pm$ 00:12:1109 : 21 : 05 ± 00 : 12 : 11
TinyImageNet 05:16:0405:16:0405 : 16 : 04
Appendix M Interval MixUp Samples
@@
Figure 9 presents representative examples of samples generated using Interval MixUp. Each row corresponds to a single mixed input (the center of a hypercube), and each column shows perturbations with increasing noise magnitude. The samples were randomly selected from the Split miniImageNet dataset. This visualization highlights how the neighborhood around each mixed point is explored, showing that small perturbations preserve the semantic content of the image, while larger ones introduce gradual variations.
@@
Figure 9: Each row corresponds to a single Interval MixUp-generated sample (hypercube center), and each column shows perturbations with increasing noise magnitude. This visualization demonstrates how added noise affects interpolated inputs, preserving semantic structure under small perturbations and gradually introducing variability.

### v3

@@ 14.3 Impact of PGD Iterations on Average Accuracy
Rysunek 7: AA of SHIELD and HNET under PGD attacks with increasing number of gradient steps, evaluated on models trained on Permuted MNIST (Fig. 7(a)), Split CIFAR-100 (Fig. 7(b)), and Split miniImageNet (Fig. 7(c)).
14.4 Impact of $\beta$ Hyperparameter Selection on Training Process of SHIELD
We evaluate the ability of SHIELD to retain knowledge across tasks, which is controlled by the hyperparameter $\beta$. In Fig. 8, we present AA curves obtained by training the model with optimal hyperparameters while varying $\beta$. The results show that $\beta$ values of $0.01$ and $1.0$ yield the best performance, ensuring stable predictions and minimal forgetting. In contrast, a small value of $\beta=0.001$ leads to significant forgetting, particularly on the 5th and 6th tasks, after learning the final task.
@@ 14.4 Impact of $\beta$ Hyperparameter Selection on Training Process of SHIELD
Rysunek 8: Comparison of AA curves for varying $\beta$ values, which control hypernetwork regularization strength in SHIELD
14.5 Performance of SHIELD on TinyImageNet dataset
We assess the scalability of SHIELD on a more challenging benchmark involving a larger number of tasks. Specifically, we use TinyImageNet, a subset of ImageNet, consisting of 200 classes, each with 500 training images and 50 validation images at a resolution of $64\times 64$ pixels. To construct the benchmark, we define 40 tasks, each composed of 5 randomly selected, non-overlapping classes. This setup allows us to evaluate SHIELD’s performance under increased task complexity and class diversity.
We compare the robustness of SHIELD with the traditional HNET approach [von2019continual] under three adversarial attacks: FGSM, PGD, and AutoAttack. Both models are trained specifically on TinyImageNet. For FGSM, we set the attack strength to $\varepsilon_{\text{attack}}=\frac{4}{255}$. For PGD and AutoAttack, we use $\varepsilon_{\text{attack}}=\frac{2}{255}$, with a PGD step size of $\delta=\frac{4}{255}$ and 100 iterations. The AutoAttack configuration is the same as used in the main experiments of the paper.
@@ 14.5 Performance of SHIELD on TinyImageNet dataset
Tabela 9: Comparison of AA after completing all tasks on the TinyImageNet dataset.
@@
From Tab. 9, we observe that even on a challenging dataset like TinyImageNet, our model achieves high AA on original (non-adversarial) test samples while also maintaining robustness against adversarial attacks. While HNET achieves competitive results, our method performs slightly better, especially under adversarial conditions.
14.6 Impact of $\kappa$ Hyperparameter Selection on Average Accuracy
To evaluate the sensitivity of our method to the hyperparameter $\kappa$, we conducted experiments on the Split CIFAR-100 dataset, measuring AA under various adversarial attacks (AutoAttack, PGD, FGSM) as well as on original samples, along with BWT. The results, averaged over 2 seeds, are presented in Tab. 10. The table compares performance across $\kappa$ values of 0.0, 0.5, and 1.0, with the best results per metric highlighted in bold.
Tabela 10: Impact of the hyperparameter $\kappa$ on AA under various attacks and backward transfer on Split CIFAR-100. Results are averaged over 2 seeds. Bold values indicate the best performance per metric.
$\kappa$ Split CIFAR-100
AutoAttack PGD FGSM Original samples
AA(%) AA(%) AA(%) AA(%) BWT
0.0 61.24 60.1 45.01 64.93 0.06
0.5 63.08 62.39 46.48 67.45 -0.41
1.0 56.58 58.94 45.54 64.55 - 8.39
From the results, $\kappa=0.5$ consistently achieves the highest average accuracy across all evaluated attacks and on original samples, demonstrating superior robustness and generalization. While BWT is slightly negative for $\kappa=0.5$, indicating minimal forgetting, it outperforms the other values, particularly $\kappa=1.0$, which exhibits significant degradation in both accuracy and BWT. In contrast, $\kappa=0.0$ shows moderate performance but falls short of the balanced improvements offered by $\kappa=0.5$. Based on these findings, we select $\kappa=0.5$ for all experiments in the main paper, as it provides the optimal trade-off between adversarial robustness, clean accuracy, and knowledge retention in continual learning settings.
15 Extended Results from Main Paper
In this section, we extend results from Table 1 to attach standard deviations. We present extended results in Table 11, 12, 13, and 14.
Tabela 11: Comparisons of AA and BWT after learning all the tasks on Permuted MNIST dataset. The results for baselines were extracted from Table 1 in [ru2024maintaining].
@@ 15 Extended Results from Main Paper
SHIELDIM $8±\pm0.0$1 97.44 $\pm$ 0.05 79.09 $\pm$ 0.87 97.96 $\pm$ 0.04 -0.05 $\pm$ 0.03
Tabela 12: Comparisons of AA and BWT after learning all the tasks on Rotated MNIST dataset. The results for baselines were extracted from Table 2 in [ru2024maintaining].
@@
Tabela 13: Comparisons of AA and BWT after learning all the tasks on Split CIFAR-100 dataset. The results for baselines were extracted from Table 3 in [ru2024maintaining].
@@
[…]
  • Note: In the original Table 3 from [ru2024maintaining], the reported accuracy for FGSM was higher than for original samples, which is inconsistent with expected adversarial behavior. We assume this was a mistake and have swapped the columns accordingly for a fair comparison.
Tabela 14: Comparisons of AA and BWT after learning all the tasks on Split miniImageNet dataset. The results for baselines were extracted from Figure 3 in [ru2024maintaining].
@@
16 Class-Incremental Learning Results
Tab. 15 presents a comparison between SHIELD and the HNET baseline in terms of adversarial robustness across four datasets: Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet, within the Class-Incremental Learning (CIL) setting. The results include AA under three adversarial attacks (AutoAttack, PGD, FGSM) as well as accuracy on original test samples, all reported after learning the full task sequence.
@@ 16 Class-Incremental Learning Results
The results reported in Tab. 15 were averaged over multiple random seeds to ensure statistical reliability. Specifically, results for Permuted MNIST, Rotated MNIST, and Split CIFAR-100 were averaged over 5 seeds, while Split miniImageNet results were averaged over 3 seeds. Due to the computational cost of AutoAttack, its results were averaged over 2 seeds for all datasets.
Tabela 15: Adversarial robustness of SHIELD across multiple datasets in the CIL scenario. We report AA under AutoAttack, PGD, and FGSM attacks, as well as on original samples after learning all tasks.
@@
6:  $\hat{t}\leftarrow\arg\min_{t}\psi_{t}$
7:  $\hat{y}\leftarrow\arg\max_{j}\hat{y}_{j}^{(\hat{t})}$
17 Comparison with AIR Method
Table 16 reports the performance of SHIELD in the AIR evaluation protocol (PGD → FGSM → None) on CIFAR-100. SHIELD achieves the highest accuracy on Task 1 and Task 3, outperforming all baselines under PGD and attaining a notably stronger final-task (clean) accuracy. Although FGSM performance is slightly below AIR, the clean accuracy of SHIELD remains substantially higher, indicating that our training does not compromise performance on original samples.
Tabela 16: Accuracy results on CIFAR-100 for the PGD $\rightarrow$ FGSM $\rightarrow$ None training sequence. Results for SHIELD are averaged over 3 random seeds. Results for baselines are extracted from Table 4 from [zhou2024defense].
Method Task 1 Task 2 Task 3
Vanilla 19.50 22.62 47.04
EWC [kirkpatrick2017overcoming] 19.88 24.48 44.83
Feature Extraction [li2017learning] 21.01 24.84 41.88
LFL [jung2016lessforgettinglearningdeepneural] 19.94 24.36 45.41
AIR [zhou2024defense] 23.12 27.04 44.16
SHIELD 23.86 $\pm$ 0.74 26.34 $\pm$ 0.43 50.72 $\pm$ 0.63
Upper bound 21.25 30.23 45.33
Since SHIELD is attack-agnostic, it cannot be directly applied as in the original AIR method. Instead, we train a single SHIELD model on the full CIFAR-100 stream and evaluate its robustness to PGD and FGSM, as well as accuracy on clean samples. For fairness, we adopt the Wide ResNet architecture used in AIR [zagoruyko2017wideresidualnetworks], but reduce it to a 10-layer variant with a widening factor of 10 to lower the number of learnable parameters and enable training on our available GPU resources. All other experimental settings follow [zhou2024defense]. We used a two-layer MLP-based hypernetwork with 100 and 50 units, respectively, and an embedding size of 512.
18 Training Algorithm
@@
19 Training Time of SHIELD
Tab. 17 reports the training time of SHIELD model across different datasets. The durations are shown in the HH:MM:SS format and include standard deviations where available. For Permuted MNIST, Rotated MNIST, Split CIFAR-100, and Split miniImageNet, results are averaged over 5 training runs using the best-performing hyperparameters. For TinyImageNet, the time is reported from a single run with no standard deviation.
Tabela 17: Comparison of training time of the SHIELD model on considered datasets.
@@ 19 Training Time of SHIELD
Permuted MNIST $00:37:59\pm 00:00:02$
Rotated MNIST $00:36:18\pm 00:00:19$
Split CIFAR-100 $02:10:23\pm 00:02:00$
Split miniImageNet $09:21:05\pm 00:12:11$
TinyImageNet $05:16:04$
20 Interval MixUp Samples
@@
Fig. 9 presents representative examples of samples generated using Interval MixUp. Each row corresponds to a single mixed input (the center of a hypercube), and each column shows perturbations with increasing noise magnitude. The samples were randomly selected from the Split miniImageNet dataset. This visualization highlights how the neighborhood around each mixed point is explored, showing that small perturbations preserve the semantic content of the image, while larger ones introduce gradual variations.
@@
Rysunek 9: Each row corresponds to a single Interval MixUp-generated sample (hypercube center), and each column shows perturbations with increasing noise magnitude. This visualization demonstrates how added noise affects interpolated inputs, preserving semantic structure under small perturbations and gradually introducing variability.
21 Theoretical Analysis of SHIELD
We now present the theoretical analysis of SHIELD together with the proof of the main theorem.
21.1 Robustness Preservation Condition Theorem
Dowód.
Using the previously introduced notation, and following Definition 3.1, our goal is to show that for any task $s$ and any $(x,y_{\text{true}})\in\mathcal{D}_{s}$, the following condition holds:
$\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h}),$ (32)
where $\boldsymbol{h}$ denotes the hypernetwork output change induced by the gradient update of the hypernetwork weights while learning the $(t+1)$-th task. This parameter change affects the classifier outputs. For each classifier logit $l_{x}$, we can write:
$f_{l_{x}}\left(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h}\right)\in\big[$ $\underline{z}^{(K)}_{l_{x}}(x;\boldsymbol{\theta}_{s,t})-\Delta_{\mathrm{max}}^{(t)}(\boldsymbol{h}),$ (33)
$\overline{z}^{(K)}_{l_{x}}(x;\boldsymbol{\theta}_{s,t})+\Delta_{\mathrm{max}}^{(t)}(\boldsymbol{h})\big].$
Since each logit changes by at most $\Delta_{\mathrm{max}}^{(t)}(\boldsymbol{h})$, it follows that
$\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})-\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})$ (34)
$\geq\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{s,t})-\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{s,t})-2\,\Delta_{\mathrm{max}}^{(s,t)}(\boldsymbol{h}).$ (35)
By the assumption stated in Eq. (14), the right-hand side of the inequality above is strictly positive. Therefore, we obtain:
$\underline{z}^{(K)}_{y_{\text{true}}}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})>\max_{j\neq y_{\text{true}}}\overline{z}^{(K)}_{j}(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h}),$ (36)
which concludes the proof. ∎
21.2 Justification via Empirical Validation of the Theoretical Condition
We now provide an empirical analysis to justify the practical efficacy of our method with respect to the theoretical conditions established in Theorem 3.1.
The theorem defines a strict, per-sample sufficient condition for preserving certified robustness: a sample $(x,y_{\text{true}})$ that is certifiably robust at time $t$ (i.e., $M(x;\boldsymbol{\theta}_{s,t})>0$) will remain so at time $t+1$ if its margin is large enough to absorb the logit change caused by the parameter update $\boldsymbol{h}$:
$M(x;\boldsymbol{\theta}_{s,t})>2\,\Delta^{(s,t)}(x,\boldsymbol{h}),$ (37)
where $\Delta^{(s,t)}(x,\boldsymbol{h})=\|f(x;\boldsymbol{\theta}_{s,t}+\boldsymbol{h})-f(x;\boldsymbol{\theta}_{s,t})\|_{\infty}$.
The Practical Proxy Argument
Explicitly enforcing this theoretical constraint for all samples at every step is computationally prohibitive in a data-free continual learning setting. We argue that the intrinsic regularization properties of our hypernetwork architecture already serve as a highly effective, automatic, and computationally efficient proxy for this hard constraint.
Our hypernetwork regularization (see Eq. (10)) restricts the change in the hypernetwork’s weights ($\boldsymbol{\Phi}$). Since the output classifier weights ($\boldsymbol{\theta}_{s,t}$) are a smooth, continuous function of $\boldsymbol{\Phi}$, this regularization effectively limits the magnitude of the resulting parameter update $\boldsymbol{h}$, consequently constraining the logit change $\Delta^{(s,t)}(x,\boldsymbol{h})$.
Empirical Validation
To validate this, Tab. 18 presents the percentage of samples that satisfy the theoretical guarantee; i.e., the percentage of samples that were initially robust (at time $t$) and for which the inequality $M(x;\boldsymbol{\theta}_{s,t})-2\,\Delta^{(s,t)}(x,\boldsymbol{h})>0$ holds after all tasks are completed. This directly measures how well our implicit regularization fulfills the explicit theorem requirements.
The results are powerful and consistent across all benchmarks. On Permuted MNIST, $97.70\%$ of samples satisfy the theoretical guarantee; on Rotated MNIST, the corresponding figure is $95.63\%$; and on Split CIFAR-100, it reaches $99.00\%$. These high rates indicate that our method’s implicit regularization effectively limits the magnitude of the logit change $\Delta^{(s,t)}(x,\boldsymbol{h})$, so that for the vast majority of inputs this change is absorbed by the margin $M(x;\boldsymbol{\theta}_{s,t})$, thereby preserving the theoretical guarantee in practice.
Tabela 18: Percentage of samples that remain robustly classified after learning the final task (Theorem 3.1). Higher values indicate better preservation of robustness. Results averaged over 5 seeds.
Task ID Permuted MNIST Rotated MNIST Split CIFAR-100
1 97.67 $\pm$ 0.42 94.49 $\pm$ 0.53 100.00 $\pm$ 0.00
2 97.09 $\pm$ 0.47 95.26 $\pm$ 0.87 95.21 $\pm$ 4.10
3 96.83 $\pm$ 0.50 94.77 $\pm$ 1.07 100.00 $\pm$ 0.00
4 97.38 $\pm$ 0.35 94.70 $\pm$ 1.53 99.04 $\pm$ 1.92
5 97.23 $\pm$ 0.47 97.07 $\pm$ 0.78 99.00 $\pm$ 1.39
6 98.02 $\pm$ 0.45 91.95 $\pm$ 4.03 98.95 $\pm$ 2.10
7 97.84 $\pm$ 0.34 95.76 $\pm$ 0.63 99.00 $\pm$ 1.17
8 98.55 $\pm$ 0.23 98.34 $\pm$ 0.14 99.76 $\pm$ 0.48
9 98.73 $\pm$ 0.59 98.33 $\pm$ 0.22 100.00 $\pm$ 0.00
Average 97.70 $\pm$ 0.11 95.63 $\pm$ 0.49 99.00 $\pm$ 0.59
In conclusion, our method successfully maintains high certified robustness across tasks. The empirical analysis confirms that the knowledge of certified robustness is automatically and sufficiently transferred from task to task via the hypernetwork’s regularization mechanism. This approach provides a practical solution that achieves the necessary condition for robustness preservation without the computational burden of explicitly enforcing the hard theoretical constraint derived from Theorem 3.1.